(a) Convert the 2-D Laplacian operator V to polar coordinates using basic 0x2 trigonometric relations between x,y,r,and p and use this result to determine an operator for 2D rotational kinetic energy about a fixed radius (b) Using the cross product formulation of the angular momentum operator below, find the components ir, ly and iz.Then show (by commutator) that lx and i, cannot be simultaneously determined but that i2 and iz can ex ey ez Tx where IM det (M) px py pz
Answer #1

Solution:

(a) 2-D laplacian to polar coordinates, the expression are shown below,

2-D equation given,

\triangledown ^2 u =\frac{\partial ^2u}{\partial x^2}+\frac{\partial ^2u}{\partial y^2}=0

Kindly post questions separately for quick answering.